Can you please help? I multiplied both side by (x-1) but i don't think it's right.
Solve for x
x+1+(x/(x-1))=1/(x-1)
Thank you

Answer by stanbon(75887)

Solve for x
x+1+(x/(x-1))=1/(x-1)
x+1 = 1/(x-1) - x/(x-1)
x+1 = [x-1-x(x-1)]/(x-1)
x+1 = [-x^2+2x-1]/(x-1)
x+1 = -[x^2-2x+1]/(x-1)
x+1 = -(x-1)^2/(x-1)
x+1 = -(x-1)
x+1 = -x+1
2x = 0
x=0
=================
Cheers,
Stan H.